# InstructAny2Pix: Flexible Visual Editing via Multimodal Instruction Following

### v2

@@ 3.4.2 Prior Alignment
To train our refinement prior, we use 2M text-image pairs from LAION-Aesthetic-V2 datasets, which consist of images with high visual quality and corresponding captions. Because the vanilla captions are of low quality, we augment the captions by recaptioning the images using BLIP2 [15]. We also make use of 2M audio-visual pairs sampled from VGG-sound [5], AudioSet [8], and SoundNet [2]. This allows the prior model to learn a broad range of textual-visual and audio-visual correspondences. We also train a version of our prior in a bidirectional way, which means our prior can also be used to predict audio tokens given images. This version is used in our data generation pipeline. The details will be discussed in the following section.
3.4.3 Instruction Tuning
We curated a diverse dataset of 500k instructions, called MM-Inst. Each instruction contains a text description of the instruction and a caption of the desired outputs. We first manually wrote a series of examples for adding, dropping, removing, or replacing objects in the scene. We also included examples of free-form audiovisual instructions, such as style adapation based on reference image or music. We then sampled 500k BLIP2-generated captions from the LAION-Aesthetic-V2 dataset and prompted a Large Language Model (LLAMA2 [30]) to generate editing instructions involving these captions, and output captions. We mark all relevant subjects such as [dog], [cat] in the instruction. During training, they were randomly replaced by the corresponding multi-modal embeddings at a fixed chance. Such embeddings are retrieved in the following way: If the text is a caption of a LAION image, we use the respective image embedding. Otherwise, we obtain the ImageBind text feature and use our prior model to randomly convert it to an audio or image feature. Additionally, we also selected 50k samples and generated pseudo ground-truth using SDXL. When there is a source image, we apply the DDIM inverse and use SDXL to generate the edited image. We visualize some pairs in Fig. 3, and examples of text pairs in Tab. 1. Our generated pairs represent a diverse range of image correspondence, including adding or removing objects, changing the foreground or background, changing the style of the image or the environment of the image. Because of known limitations of DDIM inversion, we only generated scene-level alterations and do not have examples of changing very refined details, such as the color of a person’s eyes. This is one major limitation of our data generation pipeline.
During the training, we first pretain our model using converted text embedding from the prior model as targets. We then fine-tune on 50k actual image embeddings generated by the SDXL pipeline.
Table 1: Examples from MM-Inst dataset. [.] indicates multi-modal objects marked by the large language model.
Instruction
Result
Fit the atmosphere of [a piece of music of a futuristic cityscape] to [an old black and white photo of a man, woman, and two young girls]
An image of a futuristic cityscape with the man, woman, and two young girls integrated into the scene.
Remove [the rock formation] from [the night sky is filled with stars and the Milky Way over a rock formation]
An image of the night sky filled with stars and the Milky Way.
@@ 4 Experiments
Figure 4: Multi-modal Alignment in InstructAny2Pix. We showcase results of the same instruction expressed in different modalities. We also provide comparisons with InstructPix2Pix using the same instructions. T(Ours): Text-only Instruction. MM(Ours): We replace the corresponding object with multi-modal inputs. InstructPix2Pix: We provide text-only instruction to InstructPix2Pix. Because it was trained on different styles of instructions, we reformatted the relevant instruction to make a fair comparison, e.g. “make it scary”.
Figure 5: We compare against T2I [21] on image variations. We use an official implementation of T2I based on SDXL. We select the examples from demos on the official homepage. We do not introduce novel styles. In particular, “anime,” “digital art,” and “3D rendering” are provided style choices on the HuggingFace demo. We demonstrate that taking image inputs can improve the quality of editing results, highlighting the benefits of our setup.
Figure 6: To validate the alignment of our prior, we demonstrate its any-to-image capability. We use our prior model to translate audio embedding to the image space and use our diffusion decoder to generate images. Qualitative results show that it is comparable to state-of-the-art any-to-any generation systems such as NextGPT [32]. As our work focuses on image editing, we leave more rigorous examination and extensions for future works.
Figure 7: We evaluate the effectiveness of our refinement module. Top: conditioned only on target embedding. Bottom: additionally using DDIM inverse of the original image. The leftmost column shows results with no refinement, the right three columns show results of different aesthetic scores $f$. Generally, refinement improves the result, and a higher aesthetic score leads to higher output quality.
Figure 8: Ablation Study on Control Strength. Hyper-parameters $\beta$ and $\alpha$ used in Algorithm 1 control the strength of editing. In general, a higher $\alpha$ leads to outputs that respect the spatial composition of the source image, while a higher $\beta$ leads to outputs that respect the semantics of the original image.
Figure 9: Multi-object Multi-modal Instruction Editing. Unlike previous methods that rely on an iterative chain of editing to perform complex instructions, our model can directly understand complex instructions involving multiple subjects of multiple modalities and perform them in a single turn. In each of the examples, we first provide results of iterative applying our method on two consequent instructions, then we show results of single-turn editing combining multiple instructions.
Figure 10: We provide human evaluation image-to-sound and text-to-image capabilities. For image to audio generation, we leveraged AudioLDM2 [16] to decode audio embedding to sound. The human reviewers are asked to decide which model is the most preferable among InstructAny2Pix, CoDi [28], and NextGPT [32]. We use this evaluation only to validate our data generation pipeline as we use the prior to generate audio and image embeddings.
Table 2: Alignment scores of InstructAny2Pix vs Instruct Pix2Pix on MM-Instruct-Edit dataset. Higher is better.
@@
InstructPix2Pix 0.102 0.766 0.198 3.31 3.39
InstructAny2Pix(Ours) 0.132 0.712 0.222 3.45 3.51
4.1 Multi-Modal Instruction Alignment
Due to our training scheme, our model should be able to understand the same instruction expressed in different modalities. For example, “add [sound of a dog barking]” should not yield significantly different performance than the plain text instruction “add a dog.” We validate this by comparing results of text-only instructions and multi-modal instructions. We also compare against an existing text-based instruction-edit method, InstructPix2Pix. Our model is able to follow both text-only instructions as well as multi-modal instructions. InstructAny2Pix achieves comparable or better performance compared with InstructPix2Pix.
4.2 Multi-Modal Instruction Following Capability
[…]
We also perform human evaluations. We show decisive advantages in human preference. We asked reviewers to score the alignment (Does the model follow the instructions?) and generation quality (Does the generated image look good?). We outperform InstructPix2Pix on both benchmarks. We recognized that SDXL is generally a stronger diffusion model; hence, we compared against an SDXL-based InstructPix2Pix implementation on HuggingFace. Despite both being based on SDXL, human evaluators consider our model to generate high-quality samples. We provide further discussion in the Discussion section.
4.3 Image Conditioned Style Editing
We compare against T2I [21], a popular method used to create image variations. To make it a fair comparison, we also make use of an official SDXL-based implementations. We select the image from official demos on GitHub and ask T2I to generate variations of the same image. For InstructAny2Pix, we prompt the model with an additional image reference of related styles. We demonstrated that using an image as a reference leads to higher-quality outputs.
@@ 4.4 Multi-Object Multi-Modal Instruction Editing
Unlike previous works such as [4], our methods can perform complex editing operations involving many inputs of different modalities. Fig. 9 highlights this capability. Compared with previous methods, our method drastically extends the boundary of instruction-guided image editing.
@@ 4.5 Ablations
4.5.1 Refinement Prior
To study the effectiveness of our refinement prior, we provide qualitative results of generated images in Fig. 7. We also show the effect of different aesthetic scores. As shown in the figure, the refinement process significantly improves the image quality. Additionally, we find that the aesthetic score tends to be biased towards high contrast and saturation, so depending on the use case, we can also consider reducing the conditioning score as necessary.
@@ 4.5.2 Control Strength
We visually explore the effect of hyper-parameters $\beta$ and $\alpha$ used in Algorithm 1. In general, a higher $\alpha$ leads to outputs that respect the spatial composition of the source image, while a higher $\beta$ leads to outputs that respect the semantics of the original image. When both are very high, the model simply gives the original, nearly unedited image as output. These parameters can help the user exert more refined control over desired outputs.
4.5.3 Modality Translation using Prior
Since we train our prior on paired audio-image and text-image data, it can naturally perform paired multi-modal generation. We briefly explore this ability in Eq. 2. We also qualitatively ask human evaluators to compare our work with CoDi and NextGPT. The results are shown in Fig. 10. We achieve better generation quality and are roughly comparable in alignment with these state-of-the-art models. However, this work focuses on image editing, and we use only use such experiments to validate the effectiveness of our prior in the data generation processes.
@@ 5.1 Quality of Generation
We observed that our methods generate higher quality outputs compared with previous works that also make use of SDXL [22]. We hypothesize this may be caused by our data isolation. In particular, we only train on high-quality images in a self-supervised manner, so the model does not inherit bias from editing datasets that have a lower average quality than the dataset used to pretrain SDXL. While we made this choice primarily for training efficiency, we hope future work can explore the corruption in quality of diffusion models when fine-tuned for editing.
5.2 Limitations
Our model comes with the inherited biases from pretrained diffusion and language models, in particular SDXL and Vincuna. Additionally, our editing instructions mostly consider scene-level editing because DDIM do not provide more refined control. Moreover, since it is very costly to pretrain our model end-to-end (due to the size of LLM), we cannot train on many existing instruction datasets such as [34] that incorporate Photoshop-level image editing or other refined instructions such as changing the contrast of images. Hence, it may not perform well on these benchmarks. Nevertheless, we argue that the ability to incorporate multi-modal inputs and eliminating the need for an iterative chain-of-editing for complex instructions is a major milestone in the development of instruction-based image editing. We hope future works can address these issues.
@@ 6 Conclusion
7 Acknowledgement
This research was supported by Cisco.
Supplementary Material
8 Details of Datasets
@@
We use SoundNet [2], VGG-Sound [5], and AudioSet [8] for audio-visual alignment. These datasets consist of videos with audio. We extract the audio and the middle frame from the video to create audio-image pairs. SoundNet consists of 802,724 audio-image pairs, AudioSet consists of 888,185 audio-image pairs, and VGG-Sound consists of 197,958 pairs. These numbers represent the number of valid video URLs at the time of data fetching (Oct 2023). They may differ from the original dataset size and the number of valid URLs at the time of writing. We also make use of audio captions from MusicCaps [1] and AudioCaps [13] to create text-audio pairs. These two datasets provide text captions for subsets of AudioSet. They do not introduce new audio files. We use LAION-Aesthetic-3M [26] for text-image alignment, which consists of 2,209,745 valid image URLs at the time of data fetching (Sep 2023). All these datasets are used in our prior training.

### v3

@@ 3.8 Paired Dataset
For image-text pairs, we use 2.4M text-image pairs sampled from LAION-Aesthetic-V2 datasets. Because the vanilla captions are of low quality, we augment the captions by recaptioning the images using BLIP2 [16]. For audio-text pairs, we make use of AudioCaps[13] and MusicCaps [1] dataset. For audio-image pairs, we make use of 2M audio-visual pairs sampled from VGG-sound [6], AudioSet [8], and SoundNet [3]. These dataset allow InstructAny2Pix  to learn a broad range of textual-visual and audio-visual correspondences.
3.9 Instruction Tuning Dataset
For instruction tuning, we curated a diverse dataset of 500k instructions and corresponding image pairs, called MM-Inst. The dataset generation pipeline consists of two steps: instruction generation and image generation.
In the instruction generation phase, we prompted a Large Language Model (LLAMA2 [31]) to generate creative instructions using 36 manually written examples. The example instructions include adding, dropping, removing, or replacing objects as well as other free-form instructions. To generate a diverse set of image editing instructions,we prompt the language model to create editing instructions based on sampled captions from LAION dataset. To further increase the diversity of audio-related instructions, we further prompt the language model to generate instructions involving ground truth music captions from MusicCaps and AudioCaps. We observe that without this step, the language model tends to only generate simple audio captions such as “sound of water" or “sound of rain" and fails to incorporate complex descriptions of music. In total, we generate 500k instructions. We provide further details in the appendix.
In the image generation phase, we generate corresponding image pairs of editing instructions. Unlike previous works such as InrstructPix2Pix [5] which only uses off-the-shelf caption guided image editing methods like Prompt2Prompt [11], we make use of combination of following techniques.
  • Edit image using captions and Prompt2Prompt [11]
  • Edit image using captions and Plug-and-Play [32]
  • Use DDIM[29] inversion on the source image and generate a new image using target prompt with inversed latent.
  • For object removal, we use an open-vocabulary object detector (GroundingDINO [19]) to locate the object and perform inpainting with remove the object.
  • For object addition, we first generate an image using target prompt as the target image. Then we use the detector to localize the added object and remove it through inpainting. The resulting image is used as the source image.
  • For object replacement, we first follow the removal procedure. Then we perform inpainting on the area of removed objects using the replacement object as prompts.
When the detector fails to localize the object or output low confidence scores, we fall back to the caption-based methods. Following InstructPix2Pix [5] we filter the results using CLIP scores. We additionally filter the results using the LAION aesthetic predictor and remove low quality images. We show this pipeline in Fig. 3
3.10 Evaluation Dataset
Image editing following multi-modal instruction is a novel task with no existing benchmarks. To fairly evaluate InstructAny2Pix’s performance in real-world settings, we curated 1500 manually written multi-modal instructions. We employ Amazon Mechanic Turk for this task. To generate a diverse set of prompts, we ask the MTurk workers to generate creative edit instructions using different captions sampled from LAION. We also require the MTurk workers to generate different types of edits for the same caption. We do not generate ground truth target images. For reference images and audio pieces used in the instructions, we use SDXL [24] to generate images and AudioLDM2 to generate [17] audio. Results are filtered by CLIP[25] score for images and CLAP[33] score for audios. Notable, unlike previous works which performs evaluations on samples from the same distribution of training data, our evaluation benchmark caters for diverse real-world usecases. We call this dataset MM-Inst-Test.
For image guided generation, Dreambooth[27] is a common benchmark used. However we find it inadequate to provide a holistic evaluation of the capability of multi-modal generations. Particularly, there are two limitations. Firstly, its prompts consists only two classes of live objects (cats and dogs), which in total has 9 subjects out of 30. Hence, it has very limited diversity. Secondly, its task only involve changing the background of a single subject. This setup cannot evaluate the capability of models capable of understanding multiple images. To address this gap, we propose Dreambooth++, a Dreambench dataset of more diverse prompts. It consists of 30 subject images which are evenly distributed across humans, animals, small objects, and large structures. We also includes 30 diverse background images with corresponding prompts. In total, there are 900 generation tasks. We evaluate models on this dataset using two protocols: single-image and multi-image. The single image setup is similar to Dreambooth, which requires generating a given subject under different context prompts. The multi-image task requires generating a new image by combining a subject image and a background image. We defer further details to the appendix.
Additionally, for completeness and fair comparison with existing models, we evaluate our model on 1000 samples from InstructPix2Pix dataset. Note that we do not train our model on this dataset, and the evaluation is performed in a zero-shot manner.
@@ 4 Experiments
Table 1: Multi-Modal Image Editing on MM-Inst-Test Dataset and Text-based Image Editing on InstructP2P Dataset. (I+T+A) and (T) refers to using multi-modal instruction and text-only instruction respectively.
MM-Inst InstructP2P
CLIPdir CLIPim CLIPout Win. CLIPdir CLIPim CLIPout Zero-shot
Ours(I+T+A) .099 .816 .260 -
Ours(T) .095 .856 .270 - .147 .808 .312
InstructP2P .091 .824 .243 .712 .145 .742 .241
MagicBrush .084 .807 .199 .707 .165 .760 .250
InstructDiff. .066 .940 .193 .746 .126 .857 .301
Table 2: Image Conditioned Generation on DreamBench++ Dataset. $C_{dir},C_{im},C_{out}$ is abbreviated form of CLIPdir, CLIPim and CLIPout. For multi-image setup, numbers are reported in DINOref/DINOsub format.
Single-Image Multi-Image
Cdir Cim Cout DINO Cdir Cim Cout DINO Recall
Ours(T+I) .147 .810 .260 .688 .154 .789 .309 .625/.471 .841
BLIP-Diffusion .089 .779 .231 .660 .091 .701 .292 .526/.422 .693
Kosmos-G .126 .843 .251 .683 .166 .740 .286 .485/.476 .812
Figure 5: Music-Gudied Image Variation: Music uniquely conveys emotions that are hard to describe using other modalities such as language. We recommend the reader find the corresponding music on Apple Music or other platform for optimal experience.
Figure 6: Ablation Studies of hyperparameters.
(a) Ablation Study on the effect of hyper-parameters and the refinement module. Increasing the cfg and decreasing the alpha will increase CLIPout (CLIP-T) and decrease CLIPim (CLIP-I). Adding refinement module improves both metircs.
(b) Ablation Study on Image Consistency. Notably, our method generate artifact-free results in all setup, while other methods have visiable artifact at all cfgs.
Figure 7: Multi-object Multi-modal Instruction Editing. Unlike previous methods that rely on an iterative chain of editing to perform complex instructions, our model can directly understand complex instructions involving multiple subjects of multiple modalities and perform them in a single turn. For comparison, we also show results of applying each edit sequentially.
4.1 Instruction guided Image Editing
@@
We also perform human evaluations. Human evaluators are asked to pick the preferred edit output in one-to-one comparison between $InstructAny2Pix$ and baseline methods. For a fair comparison, we use the text-only version of our method. We report win rate in Tab. 1.
We show decisive advantages in human preference and strong performance in quantitative evaluation metrics comparing with baseline methods InstructPix2Pix[5], Magicbrush [36] and InstructDiffusion [9]. Notably, we achieve competitive performance on InstructPix2Pix dataset without ever training on such dataset. This result showcases the superiority of our data generation pipeline. It incorporates a diverse range of instructions and enables our model to generalize to unseen instruction patterns.
4.2 Image Conditioned Generation
We evaluate our method on DreamBench++ dataset described in Sec. 3.10. We conduct both single-image and multi-image evaluation and compare results against Blip-Diffusion [15] and Kosmos-G [23]. In addition to metrics reported in the previous section, we additionally report DINO score which measures the image similarity. For multi-image benchmark, evaluating the DINO similarity of the entire image does not make sense as the subject is added to the scene and not necessarily occupies the entire image. To address this, we use a segmentation model [19] to segment the subject in generated images. We report the DINO similarity between the cropped subject against the referecne subject image as DINOsub, and the DINO similarity between the whole result and the background input image as DINOref. We provide quantitative metrics in Tab. 2 and qualitative comparison in the appendix. . InstructAny2Pix  shows a clear advantage in generation quality and image consistency.
@@ 4.3 Multi-Object Multi-Modal Instruction Editing
Unlike previous works such as [5], our methods is capable of performing complex editing operations involving many inputs of different modalities. Compared with previous methods, our method drastically extends the boundary of instruction-guided image editing.
@@ 4.4 Ablations
4.4.1 Refinement
To study the effectiveness of our refinement method, we compare the editing results with refinement and without refinement. We find that it leads to improvement in edit quality measured by CLIPim (+.007) and CLIPout (+.03). We also observe significant improvements in visual quality of outputs. We provide additional studies in the appendix.
@@ 4.4.2 Control Strength
An important goal of image editing is to ensure the edited images can reflect the intended changes while respecting the source image. There is usually a trade-off between these two goals. The relevant hyper parameter of InstructAny2Pix  is cfg and $\alpha$. CFG is the strength of classifier-free-guidance in the generation process. It determines the degree at which the control signal affects the generation output. We visualize the results in Fig. 6(b). We find that CFG=5 is a sweet spot for achieving high quality edit results that follows the instructions while respecting the original image. To make sure our model can follow the spatial structure of reference image, we don’t initialize the latent in the generation process from a canonical normal distribution. Instead, we sample a latent from a gaussian distribution centered at the DDIM inverse of the source image, with a variance proportional to $\alpha$. We define further details to appendix. Intuitively, $\alpha$ is the hyperparameter that trades image consistency with generation diversity. We visualize the effect of CFG and $\alpha$ in Fig. 6(a). Generally, increasing the cfg and decreasing the alpha will increase CLIPout and decrease CLIPim. For a typical use, we suggest setting $\alpha$ to 1.0.
@@ 5.1 Quality of Generation
We observed that our method generates higher quality outputs compared with previous works that also make use of SDXL [24]. We hypothesize this may be caused by our data isolation. In particular, we only train on high-quality images in a self-supervised manner, so the model does not inherit bias from editing datasets that have a lower average quality than the dataset used to pretrain SDXL. While we made this choice primarily for training efficiency, we hope future work can explore the corruption in quality of diffusion models when fine-tuned for editing.
5.2 Music Guided Generation
InstructAny2Pixuniquely enables music guided image editing, which can be quite useful in cases like automatic album generation using portraits of artists. To evaluate this potential, we perform music-guided image variation where we prompt the model with the instruction "modify [image] to convey the feeling of [music]". We find the model can successfully adapt an image to different music in subtle ways hard to be described in other modalities.
@@ 6 Conclusion
Appendix
7 Additional Experiment Results
7.1 Qualitative Comparison with Text-based Methods
We provide a qualitative comparison with methods using text instructions. We compare our results against InstructPix2Pix [5], MagicBrush [36], and Instruct Diffusion [9] in Fig. 8. For fairness, we incorporate instructions from both the InstructPix2Pix dataset and the MM-Inst-Test dataset. For MM-Inst-Test, we convert the multimodal instructions to text by replacing the multimodal token with captions of the referred image and audio. Our model shows better editing results on both datasets. This is impressive since our model is not trained on the InstructPix2Pix dataset while other methods are.
Figure 8: Qualitative Comparison Against Text-based Editing Methods. We show editing results of different models using diverse editing instructions. The top three rows are sampled from the InstructPix2Pix dataset, and the bottom three rows are sampled from the MM-Inst-Test dataset.
7.2 Qualitative Comparison with Image-conditioned Generation Methods
We provide a qualitative comparison with multi-modal generation methods that use images as references. We compare our results against BLIP-Diffusion [20] and Kosmos-G [23] in Fig. 9. Visual results show that InstructAny2Pix  outperforms these two baselines both in terms of generation quality and image consistency. We also conducted human evaluation using Amazon Mechanical Turk. We asked users to pick the best result in a side-by-side comparison of InstructAny2Pix  and baseline methods using generations of Dreambooth++ dataset. We achieved a win rate of 79.0% against BLIP-Diffusion and 86.2% against Kosmos-G.
Figure 9: Qualitative Comparison Against Image-based Generation Methods. We present generation results of different models under both single-image and multi-image setups. We employ multi-modal prompts from the Dreambooth++ dataset.
7.3 Additional Ablation Studies
7.3.1 Image Quality
We present qualitative results of multi-modal editing in Fig. 10. The results demonstrate that InstructAny2Pix  can effectively handle a diverse range of instructions involving multiple modal inputs.
Figure 10: Additional Results of Multi-Modal Editing. We showcase qualitative results of multi-modal editing. InstructAny2Pix  can handle a diverse set of instructions involving multiple modalities.
7.3.2 Image Consistency
In the main paper, we provided quantitative analysis of $\alpha$ and CFG in Sec. 4.4.2. We provide further qualitative results in Fig. 11. We compare generations at different $\alpha$ with an empty edit instruction. Recall that $\alpha$ controls the noise added to the latents obtained through DDIM inverse. When $\alpha=1$, no noise is added to inversed latents, leading to generations that closely follw the source image spatially.
Figure 11: Results of Varying $\alpha$ in the generation process. We visualize the image generation conditioned on the CLIP embedding of the source image. As $\alpha$ decreases, the generation become less consistent with the source image.
7.4 Anonymous Demo Website with Audio
Due to the nature of this work, we have created an anonymous demo website with playable audio at https://sites.google.com/view/annodemo3333/project-page. Since some music belongs to respective copyright owners, we have embedded publicly available YouTube videos instead of hosting the audio files ourselves. We are not the uploaders of these YouTube videos, and therefore they do not contain any identifiable information.
7.5 Image Quality Ablation
We sample 500 captions and evaluate the Aesthetic score and PickScore [14] for samples generated with and without the refinement process. These two metrics are known to reflect the visual quality of generated images. Aesthetic scores only consider the image quality, while PickScore additionally takes the prompt into account. On both metrics, adding the refinement module leads to considerable improvements. We show these results in Fig. 12.
Figure 12: Quantitative Ablation results of Refinement Module. We report average Aesthetic score and PickScore [14] on 500 randomly sampled captions. Aesthetic scores only consider the image quality, while PickScore additionally takes the prompt into account.
8 Additional Details about Datasets
8.1 Paired Training Data
We use SoundNet [3], VGG-Sound [6], and AudioSet [8] for image-audio pairs. These datasets consist of videos with audio. We extract the audio and the middle frame from the video to create audio-image pairs. SoundNet consists of 802,724 audio-image pairs, AudioSet consists of 2 million audio-image pairs, and VGG-Sound consists of 197,958 pairs. Particularly, out of 2 million videos in AudioSet, 1 million are under the music category. These videos can be music videos, concerts, documentaries, and other kinds of videos that use music as background, such as news programs. They naturally allow InstructAny2Pix  to learn a generic correspondence on how humans associate images and music. The rest of the audios are mostly sounds of objects. They help InstructAny2Pix  to associate the visual appearance of an object to its sound. We highlight this in Fig. 13.
@@ 8.1 Paired Training Data
We also make use of audio captions from MusicCaps [1] and AudioCaps [13] to create text-audio pairs. These two datasets provide text captions for subsets of AudioSet. They do not introduce new audio files. We use LAION-Aesthetic-3M [28] for text-image alignment, which consists of 2,209,745 valid image URLs at the time of data fetching (Sep 2023). All these datasets are used in the alignment process.
Figure 13: Example Thumbnails of Videos with Music. We show samples of thumbnails of videos from the music category of the AudioSet dataset. It incorporates diverse music-image correspondences.
8.2 MM-Inst
8.2.1 Implementation Details
In this section, we provide additional details of the generation process of multi-modal inputs. For audio and music, there are two types of inputs. The first type is generated by explicitly prompting the LLM with ground truth captions of music or sound. In this case, the corresponding audio can be directly used. The second type is created solely by the LLM. We use AudioLDM2 [18] to generate 5 samples for each of these captions and use CLAP [33] to find the sample that aligns best with the caption. For image inputs, we generate the image using SDXL from captions. For image pairs generated specifically through the segmentation+inpaint pipeline, we directly crop the images using the bounding boxes of objects and use the results as multi-modal inputs.
8.2.2 Comparison with Prior Works
Compared with the InstructPix2Pix dataset, MM-Inst has the following advantages. First, it uses BLIP-generated captions which are grounded with real images instead of the raw caption from LAION, which can be very noisy. Second, it uses a variety of techniques to generate paired data instead of solely relying on Prompt2Prompt. In particular, we observe that the segmentation+inpainting pipeline generates high-quality results for object removal when the segmentation model can correctly localize the target object. Third, we filter the data using both CLIP and Aesthetic score and consider both prompt alignment and generation quality. In contrast, InstructPix2Pix only uses CLIP score as the filtering mechanism. Lastly, MM-Inst incorporates multi-modal inputs, which makes it uniquely suitable for our multi-modal editing tasks. We showcase these differences in Tab. 3. We provide examples from both datasets in Tab. 4.
MM-Inst MM-Inst-Test InstructPix2Pix
Caption Source LAION-Aesthetics LAION-Aesthetics LAION-Aesthetics
Input Caption Generation BLIP2 BLIP2 Noisy WebData
Instruction Generation Llama 2 Human GPT-3
Paired Data Generation DDIM - Prompt2Prompt
Prompt2Prompt -
Plug-and-Play -
Segmentation+Inpaint -
Filtering Multiple Metrics Human CLIP
Modality Image,Text,Audio Image,Text,Audio Text
Size 500,000 1,500 313,010
Table 3: Comparison of MM-Inst dataset and Instruct Pix2Pix dataset.
MM-Inst
Instruction Output
Please incorporate [an image of cannon fire] into [an image of a pirate ship sailing on the high sea] An image of a pirate ship firing at a British Navy warship, fire burning on the ship
Remove [sound of car accelerating] from [an image of people driving in the countryside road] An image of a quiet countryside road
Replace [sound of dog barking] with [sound of a cute cat] for [an image of a dog at the beach] An image of a cat at the beach
Change [an image of a woman wearing sunglasses in Paris] to the style of [an image of a Renaissance painting of a noble lady] A Renaissance painting of a woman wearing sunglasses in Paris
Make [an image of a cute girl in a school uniform] fit the atmosphere of [a piece of music of stellar constellations] An image of a cute girl in a school uniform under the night sky
InstructPix2Pix
[misurina XIII… by roblfc1892] have it be a stamp Stamp… misurina XIII… by roblfc1892
[Manarola during sunrise - Cinque Terre] it is foggy Manarola during foggy sunrise - Cinque Terre
Table 4: Examples of Instructions from MM-Inst dataset and Instruct Pix2Pix dataset. Captions are marked by [.]. MM-Inst offers a better set of captions and instructions.
8.3 MM-Inst-Test
One of the challenges in generating the test dataset is that there are instances of bad formatting and low quality. We prompt each MTurk worker to generate five different edit prompts for each caption. After a batch is collected, we manually identify the problematic instructions and redistribute them to a new set of workers. In order to reach 1,500 valid instructions, we collect a total of 1795 instructions. We show the distribution of different types of instructions in table Fig. 14.
Figure 14: Distribution of Instructions in MM-Inst-Test dataset. We show the number of instructions. "Fit" refers to combining music and image, fitting an image to the style of another image, or other organic ways of combining different modalities together. "Others" include all instructions that cannot be classified as other categories, such as "transform [image] into a night scene with the sound of [sound]".
8.4 Dreambooth++
Dreambooth is commonly used to evaluate generation with image prompts; however, it only contains multiple images of the same subject, not multiple images of different objects. Moreover, the classes have limited diversity and only have two live subjects (cats and dogs). To provide a diverse and fair comparison, we propose Dreambooth++, which contains 30 subjects and 30 prompts. In total, there are 900 combinations. We use SDXL to generate all subject images. For each prompt, we also generate a background image.
We conduct evaluation on two benchmarks: The single-image setup is similar to Dreambooth, which requires generating a given subject under different context prompts. The multi-image task requires generating a new image by combining a subject image and a background image. For the multi-image task, we use a segmentation model to localize the edited objects in the new image. We report DINO similarity of the cropped subject with the reference subject image as DINOsub and the DINO similarity of the generated image with the reference background image as DINOref. We also report the recall rate of the segmentation model. These results are listed in the main paper Tab. 2.
Additionally, we provide a full list of subjects in Tab. 5 and compare it with Dreambooth. DreamBooth++ offers a more diverse range of subjects ranging from animals, humans, large structures, and small items.
Table 5: Subjects of DreamBooth++ and DreamBooth Dataset. DreamBooth++ offers a more diverse range of subjects ranging from animals, humans, large structures, and small items.
DreamBooth++ Dataset Subjects DreamBooth Subjects
a cute cat backpack
a cute dog backpack_dog
a colorful butterfly bear_plushie
a colorful bird flying low over a body of water berry_bowl
spotted horse can
an image of a squirrel candle
an african elephant walking through a grassy field cat
an image of cute anime girl cat2
an anime princess with long blonde hair and two swords clock
the woman in a black dress holding a fan colorful_sneaker
a man standing at a podium dog
a cyberpunk style dog2
an image of a scientist dog3
an image of an astronaut dog5
an artistic painting of a woman with blonde hair dog6
a wooden bridge dog7
car dog8
traffic lights duck_toy
train fancy_boot
tree grey_sloth_plushie
bicycle monster_toy
an image of a robot pink_sunglasses
tablet poop_emoji
telescope rc_car
the skull red_cartoon
vase robot_toy
wand shiny_sneaker
an image of chair teapot
an image of empty glass vase
a cowboy hat wolf_plushie